# What Matters When Repurposing Diffusion Models for General Dense Perception Tasks?

### v3

@@ Abstract
Extensive pre-training with large data is indispensable for downstream geometry and semantic visual perception tasks. Thanks to large-scale text-to-image (T2I) pretraining, recent works show promising results by simply fine-tuning T2I diffusion models for dense perception tasks. However, several crucial design decisions in this process still lack comprehensive justification, encompassing the necessity of the multi-step stochastic diffusion mechanism, training strategy, inference ensemble strategy, and fine-tuning data quality. In this work, we conduct a thorough investigation into critical factors that affect transfer efficiency and performance when using diffusion priors. Our key findings are: 1) High-quality fine-tuning data is paramount for both semantic and geometry perception tasks. 2) The stochastic nature of diffusion models has a slightly negative impact on deterministic visual perception tasks. 3) Apart from fine-tuning the diffusion model with only latent space supervision, task-specific image-level supervision is beneficial to enhance fine-grained details. These observations culminate in the development of GenPercept, an effective deterministic one-step fine-tuning paradigm tailed for dense visual perception tasks. Different from the previous multi-step methods, our paradigm has a much faster inference speed, and can be seamlessly integrated with customized perception decoders and loss functions for image-level supervision, which is critical to improving the fine-grained details of predictions. Comprehensive experiments on diverse dense visual perceptual tasks, including monocular depth estimation, surface normal estimation, image segmentation, and matting, are performed to demonstrate the remarkable adaptability and effectiveness of our proposed method. Code is at:
https://github.com/aim-uofa/GenPercept/
@@ 1 Introduction
Recent studies have explored the transferability of text-to-image (T2I) diffusion models to dense visual perception tasks, such as geometry estimation (Ke et al., 2024; Lee et al., 2024; Fu et al., 2024b; Gui et al., 2024; Ye et al., 2024), image segmentation (Van Gansbeke & De Brabandere, 2024; Lee et al., 2024), and inverse rendering (Chen et al., 2024; Kocsis et al., 2024; Zeng et al., 2024). While these works have demonstrated impressive results by repurposing diffusion models for estimating fine-grained geometric and semantic dense prediction maps, the critical design choices made in transferring diffusion models to other dense perception tasks still lack comprehensive justification. This makes it challenging to determine the optimal strategy for achieving the best performance.
@@
For example, Ke et al. (2024) align the visual perception process with the denoising process of Stable Diffusion by fine-tuning all U-Net parameters. They highlight the significance of ”multi-resolution noise” in the forward diffusion process during training, aiming to obtain clean predictions by gradually removing Gaussian noise. On the other hand, (Lee et al., 2024) modify the forward diffusion process by interpolating perception annotations with RGB images instead of using Gaussian noise, and only train the low-rank adaptation (LoRA) Hu et al. (2022) parameters while keeping the U-Net frozen. To the best of our knowledge, the effective components of these approaches have not been thoroughly investigated, leading to uncertainty about which aspects are most crucial for success.
@@
In this work, we examine the design space of repurposing diffusion models for dense visual perception tasks, and explore the key question: What factors are important when adapting diffusion models for general dense perception tasks?
@@
To answer this question, we rethink the importance of both fine-tuning protocols and fine-tuning data. From the perspective of fine-tuning protocols, we categorize recent methods into two main groups: stochastic multi-step generation and deterministic multi-step generation. We explore several critical design spaces, including the diffusion mechanism, key architectural components, training methodologies, and data quality. Our key observations are as follows: 1) The stochastic nature of diffusion models somewhat conflicts with the deterministic requirements of perception tasks, and the amount of noise used significantly impacts perception performance. 2) Strict adherence to traditional diffusion processes is unnecessary. Single-step inference provides similar performance with significantly faster execution. 3) High-quality synthetic fine-tuning data is crucial for geometry perception tasks. From the perspective of fine-tuning data quality, we conduct comprehensive dataset ablation studies on both synthetic datasets and real-world datasets.
@@
Based on the aforementioned observations, we propose GenPercept (see fig. 1), a deterministic fine-tuning paradigm featuring a remarkably simple one-step inference pipeline, an optional customized decoder, and an easily adaptable pixel-specific customized loss. We conduct extensive quantitative and qualitative experiments on a wide range of fundamental visual dense perception tasks, including monocular depth estimation, surface normal estimation, image segmentation, and matting to demonstrate the effectiveness and universality of our method.
@@
In conclusion, our contributions can be summarized as follows: 1) We systematically analyze the design space of current fine-tuning protocols, considering both model architecture and dataset selection, through comprehensive ablation studies. 2) Based on these insights, we propose GenPercept, a simple paradigm that harnesses the power of the pre-trained UNet from diffusion models for generalizable dense visual perception tasks.
@@
Standard diffusion models (Rombach et al., 2022; Chen et al., 2023; Song et al., 2020; Ho et al., 2020) define a forward process to inject Gaussian noise $\bm{\epsilon}$ into the input data and a reverse denoising process to estimate noises and generate clean images with a denoiser $\bm{v}_{\theta}$. While diffusion models generate images from noise, the visual perception tasks take RGB images $\mathbf{x}$ as input and estimate the visual perception labels $\mathbf{y}$. Aiming at transferring the prior knowledge of pre-trained diffusion models to computer vision tasks, previous works have proposed to align with the diffusion process and reformulate visual perception tasks as a multi-step denoising process, especially on the monocular depth estimation task. We review the main high-level ideas of previous algorithms in this section and offer detailed formulations in the supplementary material.
Recent works mainly explore two alignment choices. For the stochastic multi-step generation (Ke et al., 2024; Fu et al., 2024b; Gui et al., 2024), the perception tasks are reframed as an RGB image-conditioned denoising process. For training, they add random Gaussian noise $\bm{\epsilon}$ to the ground truth latent $\mathbf{z}^{(\mathbf{y})}$ in the forward diffusion process. For inference, the RGB latent $\mathbf{z}^{(\mathbf{x})}$ is concatenated with the noisy depth latent $\hat{\mathbf{z}}^{(\mathbf{y})}_{t}$ before being sent to U-Net $\bm{v}_{\theta}$ for the $t$-th denoising timestep. Clean depth latent $\hat{\mathbf{z}}^{(\mathbf{y})}_{0}$ will be available after denoising T times. The inference denoising process and forward diffusion process are shown in fig. 1 (a) and fig. 2 (a) separately. These methods perform well but face two significant drawbacks. Firstly, its stochastic nature conflicts with the deterministic nature of perception tasks, which may have a negative impact on performance. Secondly, the high-computation ensemble strategy poses challenges across various perception tasks.
To integrate the deterministic requirements of visual perception tasks with the stochastic diffusion models, deterministic multi-step generation methods such as DMP (Lee et al., 2024) reformulate the forward diffusion process as a blending interpolation between RGB images and ground-truth labels. The inference denoising process and training forward diffusion process are shown in fig. 1 (b) and fig. 2 (a), respectively. Although it addresses the uncertainty inherent in the process, the persistence of RGB texture in the generated perception predictions can detrimentally impact accuracy when visualizing the estimated outcomes. This phenomenon is particularly evident in tasks such as surface normal estimation, as exemplified in fig. 2 (c).
@@
Figure 2: Visualization of different noise forms and proportions in the forward diffusion process. (a) Compared to Gaussian Noise, interpolating RGB images with ground-truth labels can bring less image nature loss. (b) The relationship between the proportion of noise $\sqrt{\bar{\alpha}_{t}}$ and the ($\beta_{start}$, $\beta_{end}$) hyperparameters. (Please see the supplementary for detailed formulations.) (c) Low $\beta$ values during the training of deterministic multi-step generation will lead to the retention of RGB information. Increasing $\beta$ enhances the training difficulty and addresses this issue.
@@ 3 Diffusion Models for Visual Perception Tasks
In this section, we explore the necessity and highlight the findings of the multi-step stochastic diffusion mechanism, prior knowledge hides behind the architectural components, training strategy, and fine-tuning data quality. The default experimental setting is similar to Ke et al. (2024) and can be found in supplementary material. We select stochastic method Marigold (Ke et al., 2024) and the deterministic method DMP (Lee et al., 2024) as our baseline.
@@ 3.1 The Form and Proportion of Noise in the Forward Diffusion Process
In our initial attempt to integrate the core concepts of Marigold and DMP within a unified codebase, we observed that the performance of the ”baseline DMP” was significantly inferior to that of the ”baseline Marigold.” This discrepancy led us to uncover two notable phenomena: 1) Although deterministic methods are theoretically expected to outperform stochastic, multi-step generation processes for visual perception tasks due to their reduced dependency on random Gaussian noise, our findings contradicted this assumption, with deterministic methods showing inferior performance. 2) The ”multi-resolution noise” strategy introduced by Marigold substantially enhances accuracy.
To further investigate these phenomena, we hypothesize that the challenge associated with denoising from Gaussian noisy latents is substantially more difficult than that from RGB-blended noisy latents. These two different blending paradigms are illustrated in fig. 2(a). We argue that the Gaussian noisy latent less closely approximates natural images and potentially increases the complexity of training, which can enhance the training difficulty. To prove it, we modified the noise blending proportions for both methods by adjusting the $\beta$ values within the DDPM scheduler, effectively increasing the noise component during the forward diffusion process, as depicted in fig. 2(b).
Our quantitative and qualitative analyses, presented in table 1 and fig. 2(c), indicate that increasing training challenges lead to improved model performance. As the $\beta$ value increases, the impact of ”multi-resolution noise” diminishes. However, unlike DMP, which continues to show performance improvements, Marigold’s performance begins to be slightly unstable but shows a rough improvement trend when the $\beta$ value is sufficiently high. We consider it brought by the randomness of the Gaussian Noise, which dominates most components of the input latent.
When $\beta$ reaches 1, the noisy label latent input effectively becomes a pure noise latent, akin to the RGB latent in DMP. Consequently, the target output of the ”v-prediction” diffusion model shifts to the negative value of the ground truth latent, eliminating any intermediate state of blending between RGB noise latent and ground-truth latent. The denoising process becomes repeats of “predict target label $\hat{\mathbf{z}}^{(\mathbf{y})}_{0}$ from RGB image $\mathbf{z}^{(\mathbf{x})}$ and add (RGB) noise back to the RGB image $\mathbf{z}^{(\mathbf{x})}$”, whose estimation results in any timestep $t$ are exactly identical. Therefore, we propose to reduce the DDIM steps to one and call it ”deterministic one-step perception”. The resulting inference speeds are significantly faster, with performance remaining largely consistent. We consider it as “our baseline” for the subsequent analyses.
@@
Table 1: Comprehensive quantitative comparisons about the impact of noise forms and proportions in the forward diffusion process on monocular depth estimation. Visualizations of different noise forms and the effect of $\beta$ values are shown in fig. 2. The performance of DMP improves steadily, while Marigold shows initial improvements followed by a decline. When $\beta$ reaches 1, the inference process can be reduced to one step without compromising performance. “Rank” means the average rank of ten evaluation performance (smaller is better).
Type Noise Multi-res Steps ($\beta_{start}$, $\beta_{end}$) KITTI NYU ScanNet DIODE ETH3D Rank
@@
The inherent stochasticity of diffusion models can lead to minor, unstable performance fluctuations in deterministic perception tasks. By increasing the noise proportion, the multi-step generation can be simplified to a one-step fine-tuning paradigm without any loss of performance in both stochastic and deterministic methods.
@@ 3.2 Where Does the Rich Visual Knowledge Reside in Diffusion Models?
Denoiser. We reinitialize the U-Net parameter and train it from scratch on the same datasets. Without prior knowledge of large data from LAION-5B, the network performs poorly and loses the generalization. This indicates that most of the prior knowledge is stored in the denoiser module.
@@
Customized Head and Loss. The deterministic one-step perception pipeline enables customized heads and loss functions for training. By utilizing a DPT decoder or the loss functions of DepthAnythingv2 (Yang et al., 2024b), we can implement a lightweight decoder that supervises pixel-wise information at a higher resolution rather than latent features at 1/8 resolution. This approach can not only accelerate inference times but also enhance the acquisition of fine-grained details.
@@
The primary perceptual prior knowledge of diffusion models is encapsulated within the denoiser. Customized heads and loss functions can achieve enhanced inference speed and fine-grained details.

### v4

@@ Abstract
Extensive pre-training with large data is indispensable for downstream geometry and semantic visual perception tasks. Thanks to large-scale text-to-image (T2I) pretraining, recent works show promising results by simply fine-tuning T2I diffusion models for a few dense perception tasks. However, several crucial design decisions in this process still lack comprehensive justification, encompassing the necessity of the multi-step diffusion mechanism, training strategy, inference ensemble strategy, and fine-tuning data quality. In this work, we conduct a thorough investigation into critical factors that affect transfer efficiency and performance when using diffusion priors. Our key findings are: 1) High-quality fine-tuning data is paramount for both semantic and geometry perception tasks. 2) As a special case of the diffusion scheduler by setting its hyper-parameters, the multi-step generation can be simplified to a one-step fine-tuning paradigm without any loss of performance. 3) Apart from fine-tuning the diffusion model with only latent space supervision, task-specific supervision is beneficial to enhance fine-grained details. These observations culminate in the development of GenPercept, an effective deterministic one-step fine-tuning paradigm tailored for dense visual perception tasks exploiting diffusion priors. Different from the previous multi-step methods, our paradigm offers a much faster inference speed, and can be seamlessly integrated with customized perception decoders and loss functions for task-specific supervision, which can be critical for improving the fine-grained details of predictions. Comprehensive experiments on a diverse set of dense visual perceptual tasks, including monocular depth estimation, surface normal estimation, image segmentation, and matting, are performed to demonstrate the remarkable adaptability and effectiveness of our proposed method.
@@ 1 Introduction
Recent studies have explored the transferability of text-to-image (T2I) diffusion models to dense visual perception tasks, such as geometry estimation (Ke et al., 2024; Lee et al., 2024; Fu et al., 2024b; Gui et al., 2024; Ye et al., 2024), image segmentation (Van Gansbeke & De Brabandere, 2024; Lee et al., 2024), and inverse rendering (Chen et al., 2024; Kocsis et al., 2024; Zeng et al., 2024). While these works have demonstrated impressive results by repurposing diffusion models for estimating geometric and semantic dense prediction maps, the critical design choices made in transferring diffusion models to other dense perception tasks still lack comprehensive justification. This makes it challenging to determine the optimal strategy for achieving optimal performance.
@@
For example, Ke et al. (2024) align the visual perception process with the denoising process of Stable Diffusion by fine-tuning all U-Net parameters. They highlight the significance of “multi-resolution noise” in the forward diffusion process during training, aiming to obtain clean predictions by gradually removing Gaussian noise. On the other hand, Lee et al. (2024) modify the forward diffusion process by interpolating perception annotations with RGB images instead of using Gaussian noise, and only train the low-rank adaptation (LoRA) (Hu et al., 2022) parameters while keeping the U-Net frozen. To our knowledge, the effective components of these approaches have not been thoroughly investigated, and it is unclear which design choices contribute most to the success.
@@
In this work, we examine the design space of repurposing diffusion models for dense visual perception tasks, and attempt to answer the key question: What are the important design choices when adapting diffusion models for general dense perception tasks?
@@
To answer this question, we rethink the importance of both fine-tuning protocols and fine-tuning data. From the perspective of fine-tuning protocols, we categorize recent methods into two main groups: stochastic multi-step generation and deterministic multi-step generation. We explore several critical design choices, including the diffusion mechanism, key architectural components, training methodologies, and data quality. Our key observations are as follows: 1) By setting the hyperparameters of the diffusion scheduler to particular values, the multi-step generation can be simplified to a one-step fine-tuning paradigm without any loss of performance. 2) Strict adherence to traditional diffusion processes appears to be unnecessary. Single-step inference provides similar performance with significantly faster execution. 3) High-quality synthetic fine-tuning data is crucial for several perception tasks. From the perspective of fine-tuning data quality, we conduct comprehensive dataset ablation studies on both synthetic datasets and real-world datasets.
@@
Based on the aforementioned observations, we propose GenPercept (see Fig. 1), a deterministic fine-tuning paradigm featuring a remarkably simple one-step inference pipeline, an optional customized decoder, and an easily adaptable pixel-wise loss. We conduct extensive quantitative and qualitative experiments on a wide range of visual dense perception tasks, including monocular depth estimation, surface normal estimation, image segmentation, and matting to demonstrate the effectiveness and generalization capability of our method.
@@
In conclusion, our contributions can be summarized as follows: 1) We systematically analyze the design space of fine-tuning protocols, considering both model architecture and dataset selection, through comprehensive ablation studies. 2) Based on these insights, we propose GenPercept, a simple paradigm that harnesses the power of the pre-trained UNet from diffusion models for generalizable dense visual perception tasks.
@@
We take the latent diffusion model as an example. To model the data distribution, the idea of the diffusion model (Rombach et al., 2022; Chen et al., 2023; Song et al., 2020; Ho et al., 2020) is to randomly sample a noise $\mathbf{z}^{(\mathbf{y})}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and sequentially denoise it into a $\mathbf{z}^{(\mathbf{y})}_{0}$, which is distributed according to the data. In the forward diffusion process, $\mathbf{z}^{(\mathbf{y})}_{t}$ is sampled by $\mathbf{z}^{(\mathbf{y})}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}^{(\mathbf{y})}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$, where $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and $\bar{\alpha}_{t}=\prod_{s=1}^{t}{(1-\beta_{s})}$. The variance schedule $\{\beta_{t}\in(0,1)\}_{t=1}^{T}$ is interpolated between $\beta_{start}$ and $\beta_{end}$ with $T$ steps, where larger values of $(\beta_{start},\beta_{end})$ correspond to smaller $\bar{\alpha}_{t}$ values, i.e., smaller proportion of noise. In the reverse process, Salimans & Ho (2021) use a “v-prediction” objective, where a denoiser $\bm{v}_{\theta}$ minimizes the following:
To fully leverage the pre-trained prior of the diffusion models for dense prediction tasks, previous works have reformulated these tasks as a multi-step denoising process, especially on monocular depth estimation. Given a data pair ($\mathbf{z}^{(\mathbf{x})}$, $\mathbf{z}^{(\mathbf{y})}$) where $\mathbf{z}^{(\mathbf{x})}$ is the observation and $\mathbf{z}^{(\mathbf{y})}$ is the prediction target, stochastic multi-step generation methods (Ke et al., 2024; Fu et al., 2024b; Gui et al., 2024) such as Marigold (Ke et al., 2024) add $\mathbf{z}^{(\mathbf{x})}$ as an additional input to the denoiser $\bm{v}_{\theta}$, and use $\bm{v}_{\theta}(\mathbf{z}^{(\mathbf{y})}_{t},\mathbf{z}^{(\mathbf{x})},t)$ to predict $\mathbf{z}^{(\mathbf{y})}$. By contrast, deterministic multi-step generation methods such as DMP (Lee et al., 2024) take the observation $\mathbf{z}^{(\mathbf{x})}$ as a deterministic noise and compose $\mathbf{z}^{(\mathbf{y})}_{t}$ as a blend between $\mathbf{z}^{(\mathbf{x})}$ and $\mathbf{z}^{(\mathbf{y})}$:
The denoising process and forward diffusion process of these two categories are illustrated in Fig. 1 and Fig. 2 (a). We offer detailed formulations in the supplementary material.
@@
Figure 2: Illustration of different noise forms and proportions in the forward diffusion process. (a) Visualization of interpolating ground-truth labels with Gaussian noise and RGB noise. (b) The relationship between the noise proportion $\sqrt{\bar{\alpha}_{t}}$ and the ($\beta_{start}$, $\beta_{end}$) hyperparameters. (c) Small $(\beta_{start},\beta_{end})$ values during the training of deterministic multi-step generation tend to lead to unclean estimation, which contains some RGB information. Enlarging them may alleviate this issue.
@@ 3 Diffusion Models for Visual Perception Tasks
In this section, we explore the necessity and highlight the findings of the multi-step diffusion mechanism, the architectural components, training strategy, and fine-tuning data quality. We select the stochastic method Marigold (Ke et al., 2024) and the deterministic method DMP (Lee et al., 2024) as our baseline methods. The default experimental setting here is similar to Ke et al. (2024) and can be found in the supplementary material.
@@ 3.1 The Form and Proportion of Noise in the Forward Diffusion Process
For the training process of Marigold and DMP, the timestep $t$ is sampled to control the proportion of noise added to the ground truth, and the network is trained to recover a clean ground truth from a noisy latent. For smaller timesteps such as “$t=200$”, as illustrated in Fig. 2(a), the input to the network retains significant ground truth information (e.g., the purple color of the normal). Therefore, we hypothesize that a certain level of ground-truth label information being part of the input makes the network comparatively be easier to recover the clean ground truth latent than the case of the absence of any ground truth information during training. This can limit the network’s capacity to learn comprehensive knowledge and lead to unsatisfactory performance, as it is known that networks can become lazy that tend to exploit “shortcuts”—the ground truth labels in the input in our case. On the other hand, for inference there is no such ground truth available, causing disparity of input signals between inference and training. We note that T2I tasks may suffer less due to their stochastic nature, namely, converting a text prompt to a generated image is one-to-many mapping process.
To alleviate this issue, we attempt to control the blending proportion by changing the ($\beta_{start}$, $\beta_{end}$) values of the diffusion model’s DDPM scheduler. As shown in Fig. 2(b) and Fig. 2 of the supplementary, training with a ($\beta_{start}$, $\beta_{end}$) value of (1.0, 1.0) achieves the best rank performance for both Gaussian noise and RGB noise, which is demonstrated in Table 1. Rather than achieving consistent performance improvement while increasing the noise proportion, we observed that Marigold’s performance begins to be slightly unstable when the ($\beta_{start}$, $\beta_{end}$) values are sufficiently high. Experiments of varying the random seed during both the training and inference process are conducted to rule out the influence of randomness. The results show that we may not be able to exactly find a set of unique values for ($\beta_{start}$, $\beta_{end}$) to achieve the best accuracy, as the final accuracy can be affected by a few other factors besides the aforementioned one.
Additionally, when ($\beta_{start}$, $\beta_{end}$) are equal to 1, the noise proportion $\sqrt{\bar{\alpha}_{t}}$ is equal to 0, and the formulation of DMP can be derived from Eq. 1 and Eq. 8 as follows.
$\begin{gathered}\mathbf{z}_{t}=\mathbf{z}^{(\mathbf{y})}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}^{(\mathbf{y})}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{z}^{(\mathbf{x})}=\mathbf{z}^{(\mathbf{x})},~{}~{}t=[1,...,T],\\ \mathcal{L}=\mathbb{E}_{\mathbf{z}^{(\mathbf{y})},\bm{\epsilon}\sim\mathcal{N}(0,I),t\sim\mathcal{U}(T)}\left\|-\mathbf{z}^{(\mathbf{y})}-\bm{v}_{\theta}(\mathbf{z}_{t},t)\right\|^{2}_{2}.\end{gathered}$ (3)
In this case, the output of the denoiser $\bm{v}_{\theta}(\cdot,\cdot)$ is enforced to learn the negative value of the ground truth latent for each step, and the multi-step denoising is equivalent to the single-step denoising. We propose to reduce the DDIM steps of DMP to one and call it “deterministic one-step perception”. The resulting inference can be significantly faster, with performance remaining almost unchanged. We name this as “our baseline” for the subsequent analysis.
@@
Table 1: Comprehensive quantitative comparisons about the impact of noise forms and proportions in the forward diffusion process on monocular depth estimation. Visualizations of different noise forms and the effect of ($\beta_{start}$, $\beta_{end}$) values are shown in Fig. 2. The performance of DMP improves steadily, while Marigold shows initial improvements followed by a decline. When $\beta_{start}$ and $\beta_{end}$ are equal to 1, the inference process can be reduced to one step without compromising performance. “Rank” means the average rank of ten evaluation performance (smaller is better).
[…]
@@
By setting the ($\beta_{start}$, ${\beta_{end}}$) values to 1, the multi-step generation is simplified to a one-step fine-tuning paradigm without any loss of performance in both stochastic and deterministic methods, e.g., Marigold (Ke et al., 2024) and DMP (Lee et al., 2024) respectively.
@@ 3.2 Where Does the Rich Visual Knowledge Reside in Diffusion Models?
Denoiser. We reinitialize the U-Net parameters and train the network from scratch on the same datasets. Without prior knowledge of large data from LAION-5B, the network performs poorly and loses the generalization capability. This indicates that most of the prior knowledge is stored in the U-Net module.
@@
Customized Head and Loss. The deterministic one-step perception pipeline enables customized heads and loss functions. By utilizing a DPT decoder (Ranftl et al., 2021) and the loss functions of DepthAnythingv2 (Yang et al., 2024b), we can implement a lightweight decoder that supervises pixel-wise information at a higher resolution rather than latent features at 1///8 resolution. This approach can accelerate inference times and enhance the acquisition of fine-grained details.
@@
The primary perceptual prior knowledge of diffusion models is encapsulated within the U-Net of the diffusion model. Customized heads and loss functions offers flexibility and may lead to faster inference speed and improved results.